OpenStudy (anonymous):

Which ordered pair is NOT a solution of the inequality, y < 3 - 2x? A. (-1,3) B. (-3, -10.5) C. (0, -6) D. (2, -3)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (texaschic101):

lets sub and find out.. y < 3 - 2x y < 3 - 2x (-1,3)...x = -1 and y = 3 (0,-6)...x = 0 and y = -6 now sub sub 3 < 3 - 2(-1) -6 < 3 - 2(0) 3 < 3 + 2 -6 < 3 - 0 3 < 5....this is true -6 < 3....this is true y < 3 - 2x y < 3 - 2x (-3,-10.5)...x = -3 and y = -10.5 (2,-3)...x = 2 and y = -3 sub sub -10.5 < 3 - 2(-3) -3 < 3 - 2(-3) -10.5 < 3 + 6 -3 < 3 + 6 -10.5 < 9......this is true -3 < 9...this is true they are all correct....Is the problem written correctly ?

OpenStudy (anonymous):

No I'm sorry! It's actually A. (-1,3) B. (6,-7) C. (0,4) D.(-4, 5)